the velocity of a body becomes(1/nTH)of its velocity while passing through a plank,then the no of such planks required to just stop the bullet -------

Grade:

1 Answers

vikas askiitian expert
13 years ago

 by using v^2=u^2+2as

 final velocity becomes u/n after passing through plank

 let width of each plank be d , thus from here we can find the retardation a=(u^2)-(u^2/n^2) /2d

  if N number of planks are placed, after passing through Nth plank the velocity of bullet becomes zero

 putting v=0 and s=Nd

we get 

              N=(n^2)/(n^2-1)